OpenStudy (anonymous):

new here. got a ? tho i need help finding solutions to 2sinx=1-sinx

OpenStudy (amistre64):

2sin(x) = 1 - sin(x) 2sin(x) + sin(x) = 1 sin(x)(2+1) = 1 sin(x) = 1/3 x = sin^1(1/3)

OpenStudy (anonymous):

sweet that nailed it down